### v2

# Co-rewarding: Stable Self-supervised RL for Eliciting Reasoning in Large Language Models

Abstract
While reinforcement learning with verifiable rewards (RLVR) is effective to improve the reasoning ability of large language models (LLMs), its reliance on human-annotated labels leads to the scaling up dilemma, especially for complex tasks. Recent self-rewarding methods investigate a label-free alternative to unlock the reasoning capabilities of LLMs, yet they frequently encounter the non-negligible training collapse issue, as the single-view supervision signal easily forms the self-consistent illusion, yielding the reward hacking. Inspired by the success of self-supervised learning, we propose Co-rewarding, a novel self-supervised RL framework that improves training stability by seeking complementary supervision from another views. Specifically, we instantiate Co-rewarding in two ways: (1) Co-rewarding-I is a data-side instantiation that derives reward signals from contrastive agreement across semantically analogous questions; and (2) Co-rewarding-II is a model-side instantiation that maintains a slowly-updated reference teacher with pseudo labels to realize self-distillation. Intuitively, such instantiations introduce different levels of discrepancy to increase the difficulty of training collapse on trivial reasoning solutions. Empirically, Co-rewarding exhibits stable training across various setups, and outperforms other self-rewarding baselines by $3.31\%$ improvements on average on multiple mathematical reasoning benchmarks, especially by $7.49\%$ on Llama-3.2-3B-Instruct. Notably, Co-rewarding reaches or even surpasses RLVR with ground-truth (GT) label in several cases, such as a Pass@$1$ of $94.01\%$ on GSM8K with Qwen3-8B-Base remarkably higher than GT. Our code is publicly available at https://github.com/tmlr-group/Co-rewarding.
Figure 1: Performance overview. Reasoning comparison of Pass@1 value and validation curves. Our Co-rewarding achieves better and more stable (without collapse) training than other baselines.
1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024; Qwen et al., 2025) have demonstrated remarkable general-purpose capabilities in a wide range of linguistic tasks (Hendrycks et al., ). To further elicit their reasoning ability in complex scenarios, reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; Yu et al., 2025) is developed for post-training with externally verifiable signals like program execution results (Luo et al., 2025) or mathematical equivalence (Shao et al., 2024). Despite the impressive improvement, the reliance on high-quality ground-truth (GT) labels of RLVR remains as a major bottleneck (Ouyang et al., 2022; Bai et al., 2022) in the spirit of the scaling law, which subsequently motivates the emerging exploration of self-rewarding methods with unlabeled data (Zhao et al., 2025b; Zuo et al., 2025; Zhang et al., 2025b).
One prominent line of such label-free methods leverages the internal signals (e.g., entropy (Zhang et al., 2025c; Prabhudesai et al., 2025) and self-certainty (Zhao et al., 2025b)) to strengthen the confidence of the model in reasoning. Another critical line seeks the answer-level consensus (Zuo et al., 2025; Shafayat et al., 2025) to construct pseudo labels as reward basis. while effective initially, these self-rewarding approaches frequently exhibit non-negligible training collapse (Zhang et al., 2025d) (indicated as right of Figure 1), which limits the scalability of such label-free training manners.
The collapse phenomenon stems from reward hacking (Laidlaw et al., 2025) under self-consistent illusion: the reward signal is internally produced by the policy model from a single-view data perspective, which is easily trapped by trivial solutions along with training (see Figure 7). Specifically, for entropy- or certainty-based rewards, the policy model may concentrate probability mass on a small set of tokens and produce repetitive strings that minimize entropy or maximize self-certainty (Zhang et al., 2025d). And for consensus-based rewards, the policy model can converge to a consistent yet incorrect answer that attains high consensus across rollouts (Shafayat et al., 2025). Overall, the policy model continually reduces uncertainty without sustained gains in correctness, inflating the reward but eroding exploration and diversity. It ultimately collapses when a persistent hacking strategy emerges.
To this end, we introduce Co-rewarding, a self-supervised RL framework that seeks complementary supervision from another views, inspired by self-supervised learning (Chen et al., 2020; Grill et al., 2020; Caron et al., 2021). Conceptually, one fundamental characteristic of self-rewarding methods lies on that supervision intertwined with current policy on single-view outputs, for which we propose to seek reasoning invariance across different views (see Figure 2). Specifically, we investigate two initiations of Co-rewarding: (1) Co-rewarding-I: a data-side initiation that constructs rewards via contrastive agreement across semantically analogous questions, each providing pseudo labels for the other; and (2) Co-rewarding-II: a model-side initiation that introduces an extra teacher with dynamically updated policy and provides stable pseudo-labels insulated from current online policy.
By introducing cross-view supervision on data and decoupling the reward signal from the current policy, Co-rewarding effectively mitigates training collapse and yields stable self-supervised RL training. Extensive experiments across multiple datasets validate the stability and superiority of Co-rewarding, compared to several recent baselines across several LLM families including Qwen3/2.5 and Llama. Notably, both Co-rewarding-I and -II reach or exceed training with ground-truth labels in several settings, such as achieving up to $94.01\%$ Pass@1 on GSM8K. Our main contributions are
  • We introduce a new perspective, from self-supervised learning, to elicit reasoning capability via another views of supervision, which prevents the model from training collapse (Section 3.1).
  • We propose Co-rewarding, a novel self-supervised RL framework that is initiated by the data and model sides to construct self-generate rewards to promote stably reasoning elicitation (Section 3.2).
  • We empirically demonstrate the general effectiveness of Co-rewarding to achieve superior reasoning performance on LLMs, and also present various ablation studies and further analyses (Section 4).
2 Preliminary
Problem Setups. Given a LLM $\pi_{\theta}$ parameterized by $\theta$ and a dataset $\mathcal{D}$ of question–answer pairs $(x,a)$, the model generates a response $y\sim\pi_{\theta}(\cdot\mid x)$ autoregressively. Let $y=(y_{1},\dots,y_{n})$, where each token is sampled as $y_{t}\sim\pi_{\theta}(\cdot\mid x,y_{<t})$ given the generated prefix $y_{<t}$. We consider the LLM outputs a stepbystep reasoning trace and a final answer. A verifiable reward function $r(a,y)$ compares the extracted answer $\mathrm{ans}(y)$ with the ground truth $a$ as follows:
$r(a,y)=\left\{\begin{aligned} &\text{1}&&\text{If}~\mathrm{ans}(y)~\text{is correct with answer}~a,\\ &\text{0}&&\text{If}~\mathrm{ans}(y)~\text{is incorrect with answer}~a.\end{aligned}\right$. (1)
Then, the general objective of training LLM for reasoning via RLVR (Shao et al., 2024; Yu et al., 2025) can be formulated with the policy model $\pi_{\theta}$ as follows:
$\max_{\pi_{\theta}}\mathbb{E}_{(x,a)\in\mathcal{D},~y\sim\pi_{\theta}(x)}[r(a,y)-\beta\cdot\mathrm{KL}[\pi_{\theta}(y|x)||\pi_{\textrm{ref}}(y|x)]],$ (2)
where $\pi_{\textrm{ref}}$ is an initial reference policy, and $\beta$ is a coefficient controlling the KL divergence to prevent excessive deviation from the reference model. Intuitively, the training target is to maximize the reward in passing specific reasoning questions while maintaining the general capability of LLM.
Group Relative Policy Optimization (GRPO). In practice, we adopt GRPO (Shao et al., 2024), a widely used and representative optimization method for objective Eq. (2) that estimates the advantage by normalizing the reward across multiple sampled outputs for the same question. Specifically, for a given question $x$, GRPO samples $G$ outputs from the old policy $\pi_{\textrm{old}}$ as $\{y_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|x)$. It then computes a reward for each output $y_{i}$ via a deterministic reward function, forming a group of rewards $\{r(a,y_{i})\}_{i=1}^{G}$ to estimate the advantage $\hat{A}_{i}$ as follows:
$\hat{A}_{i}=\frac{r(a,y_{i})-\text{mean}(\{r(a,y_{i})\}_{i=1}^{G})}{\text{std}(\{r(a,y_{i})\}_{i=1}^{G})}.$ (3)
Then, the target policy is optimized by maximizing the advantage while ensuring the policy model remains close to the reference policy:
[…]
where
$c_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}|x,y_{i,<t})}{\pi_{\theta_{\textrm{old}}}(y_{i,t}|x,y_{i,<t})},\;\mathbb{D}_{\textrm{KL}}(\pi_{\theta}||\pi_{\textrm{ref}})=\frac{\pi_{\theta}(y_{i,t}|x,y_{i,<t})}{\pi_{\textrm{ref}}(y_{i,t}|x,y_{i,<t})}-\log\frac{\pi_{\textrm{ref}}(y_{i,t}|x,y_{i,<t})}{\pi_{\theta}(y_{i,t}|x,y_{i,<t})}-1.$ (5)
[…]
Figure 2: Illustration of Co-rewarding framework: Unlike single-view methods that rely only on internal reward signal on original question (a), Co-rewarding introduces complementary supervision. On the data side (b), paraphrased questions yield pseudo-labels for cross-reference. On the model side (c), teacher model isolated from current policy provides stabilized pseudo-labels for updates.
3 Co-rewarding
In the following, we present Co-rewarding in detail, a novel self-supervised RL framework for LLM to elicit the latent reasoning capability through the intuition of seeking complementary supervision.
3.1 Conceptual Philosophy: invariance beyond the single-view
At the core of self-rewarding methods lies a fundamental tension: the model derives supervisory signals from its own outputs, inevitably intertwining supervision with policy and risks collapse. True reasoning competence, however, cannot be reduced to the mere correctness of isolated answers. It should instead reflect invariance that extends beyond the single-view output for consistency. This calls for training signals that remain valid across different data views or persist throughout the temporal evolution of the model, providing a more reliable basis on which self-supervised RL can rely. In this aspect, stability arises from invariance that prevents reasoning against superficial variations in data and guides the model towards increasingly valid reasoning trajectories throughout training.
This philosophy yields our Co-rewarding framework, whose core idea is to ground self-supervised RL in invariance rather than the suspicious single-view feedback. We instantiate it in two orthogonal ways: by enforcing analogy-invariance on the data side (Co-rewarding-I) and by disentangling supervision through temporal invariance on the model side (Co-rewarding-II).
3.2 Two Initiations of Co-rewarding Framework
Co-rewarding-I: on the Data Side. Inspired by contrastive learning, such as SimCLR (Chen et al., 2020) and InfoNCE (Oord et al., 2018), where two views of the same data are encouraged to have similar representations, we hypothesize an analogy-invariance inductive property of LLMs in eliciting reasoning capacity: questions that share the same mathematical essence but differ in surface form (e.g., via paraphrasing, background substitution, or reformatting) should elicit the comparably valid and similar reasoning results. This forms the foundation for a self-referential training signal: contrastive agreement among different question variants can serve as an optimization proxy. Co-rewarding-I defines contrastive agreement as a principle that aligns model reasoning outputs, treating consistent inter-view agreement as a signal for valid inference. This complements single-view self-rewarding strategies by introducing a form of collective validity verification with broader input consideration.
Building upon the discussed contrastive agreement, we initiate our Co-rewarding-I as illustrated in Figure 2. Formally, its learning objective can be formulated as follows based on GRPO:
$\begin{split}\mathcal{J}_{\textrm{Co-rewarding-I}}(\theta)\;=\;\underbrace{\mathbb{E}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x\in\mathcal{D}},\{{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}}}\}_{i=1}^{G}\sim\pi_{\theta_{\textrm{old}}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\cdot|x})}\mathcal{R}_{\theta}(\hat{A})}_{\mathcal{J}_{\textrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}original}}}(\theta)}\;+\;\underbrace{\mathbb{E}_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}x^{\prime}\in\mathcal{D^{\prime}}},\{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}{y_{i}}^{\prime}}\}_{i=1}^{G}\sim\pi_{\theta_{\textrm{old}}}({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\cdot|x^{\prime}})}\mathcal{R}_{\theta}(\hat{A}^{\prime})}_{\mathcal{J}_{\textrm{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}rephrased}}}(\theta)},\end{split}$ (6)
where the relative advantages are estimated by the cross-refereed supervision as follows:
$\hat{A}_{i}=\frac{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y^{\prime}_{\textrm{v}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})-\text{mean}(\{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y^{\prime}_{\textrm{v}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})\}_{i=1}^{G})}{\text{std}(\{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y^{\prime}_{\textrm{v}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})\}_{i=1}^{G})},\;\hat{A}^{\prime}_{i}=\frac{r({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{\textrm{v}}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y_{i}^{\prime}})-\text{mean}(\{r({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{\textrm{v}}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y_{i}^{\prime}})\}_{i=1}^{G})}{\text{std}(\{r({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{\textrm{v}}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y_{i}^{\prime}})\}_{i=1}^{G})}.$ (7)
Specifically, given a set of original questions, we utilize the rephrased version that keeps the semantical equivalence for the model to respond, and then collect the self-generated pseudo-labels based on the majority voting mechanism (Shafayat et al., 2025) as follows to supervise learning on the counterparts,
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{\textrm{v}}}\leftarrow{\arg\max_{y*}\sum_{i=1}^{G}\mathrm{1}[\mathrm{ans}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})=\mathrm{ans}(y*)]},\quad{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y_{\textrm{v}}^{\prime}}\leftarrow{\arg\max_{y*}\sum_{i=1}^{G}\mathrm{1}[\mathrm{ans}({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y_{i}^{\prime}})=\mathrm{ans}(y*)]}.$ (8)
The overall pipeline can be viewed as a dual-path structure with cross-reference in the reward shaping process, it may also be compatible with other self-generated feedbacks (Wang et al., 2022) on the output-side information due to the generality of the core idea. While in the current version, we choose the majority voting mechanism in the implementation for the empirical effectiveness and simplicity.
We summarize the pseudo code of Co-rewarding-I in Algorithm 1. Our contrastive objective operates on self-generated reasoning answers, encouraging the model to align its reasoning results to different questions that share the similar semantic intent. Formally, for each input question, the signal of Co-rewarding-I increases when the model’s output is consistent with the majority answer obtained from its analogical counterparts, and decreases when it diverges. This contrastive agreement promotes semantic invariance, implicitly increasing the difficulty of reaching trivial solutions to obtain the reward (e.g., achieving the arbitrary answers but consistent on single input) by involving data-side analogy. We leave a more intuitive case study in the Appendix D.9 to present the rephrased questions.
Co-rewarding-II: on the Model Side. On the data side, our Co-rewarding-I provides complementary supervision by involving question analogy, while its pseudo-labels are still generated by the current online policy and may depend on rephrasing quality; consequently, supervision remains partially entangled with the policy. Inspired by self- or weakly supervised methods like the representative BYOL (Grill et al., 2020), DINO (Caron et al., 2021), and Co-teaching (Han et al., 2018), which share the common intuition of introducing an auxiliary network to provide supervision beyond the current model, we initiate Co-rewarding-II from another view of complementary supervision: a model-side strategy that sources pseudo-labels from a teacher reference, which disentangle the self-supervision reward from the online policy. To avoid the heavy cost of adding and maintaining another LLM in training, Co-rewarding-II reuses the GRPO reference model as the teacher to generate the rollouts and produce pseudo-labels. In particular, the teacher is dynamically updated as an exponential moving average (EMA) of the student policy to ensure pseudo-label quality improving as the policy improves.
Intuitively, we illustrate Co-rewarding-II in Figure 2. Its learning objective can be formulated as:
$\mathcal{J}_{\textrm{Co-rewarding-II}}^{(k)}(\theta)\;=\;\mathbb{E}_{x\in\mathcal{D},\underbrace{\{y_{i}\}_{i=1}^{G}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pi_{\theta_{\textrm{old}}}^{(k)}}(\cdot|x),}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{policy student rollout}}}\underbrace{\{\tilde{y}_{j}\}_{j=1}^{\tilde{G}}\sim{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{\pi}_{\textrm{ref}}^{(k)}}(\cdot|x)}_{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\text{reference teacher rollout}}}}\mathcal{R}_{\theta}(\hat{A}^{(k)}),$ (9)
where $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}}\}_{i=1}^{G}$ are policy rollouts and $\{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{j}}\}_{j=1}^{\tilde{G}}$ are reference teacher rollouts at the $k$-th training step, and the estimated advantage $\mathcal{R}(\hat{A}^{(k)})$ is computed as follows:
$\hat{A}_{i}^{(k)}=\frac{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{(k)}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})-\textrm{mean}(\{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{(k)}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})\}_{i=1}^{G})}{\textrm{std}(\{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{(k)}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})\}_{i=1}^{G})},\;{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{(k)}}=\arg\max_{y*}\sum_{j=1}^{\tilde{G}}\mathbf{1}[\textrm{ans}({{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{j}}})=\textrm{ans}(y*)],$ (10)
where the pseudo label $\tilde{y}_{\textrm{v}}^{(k)}$ is obtained via majority voting from reference rollouts, and the reference model is updated via an EMA with the policy to play a role of a slowly updated teacher:
${\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{\pi}_{\textrm{ref}}^{(k)}}\leftarrow\alpha^{(k)}\cdot{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{\pi}_{\textrm{ref}}^{(k-1)}}+(1-\alpha^{(k)})\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pi_{\theta_{\textrm{old}}}^{(k)}},\;\alpha^{(k)}=1-\frac{(\alpha_{\textrm{end}}-\alpha_{\textrm{start}})}{2}\left(1+\cos\left(\frac{\pi k}{K}\right)\right)$ (11)
where $\alpha^{(k)}\in(0,1)$ is the EMA weight, updated according to a cosine annealing schedule from $\alpha_{\textrm{start}}$ to $\alpha_{\textrm{end}}$, such that the teacher is updated rapidly at the beginning and progressively more slowly, thereby evolving smoothly and remaining temporally decoupled from the current online policy.
We summarize the pseudo code of Co-rewarding-II in Algorithm 2. This design can be interpreted as a kind of self-distillation, in which a slowly updated teacher supervises a faster-moving student. Such a paradigm breaks the single-step on-policy feedback loop inherent in existing self-rewarding methods (Zhao et al., 2025b; Prabhudesai et al., 2025; Shafayat et al., 2025), raises the cost of exploiting trivial low-entropy shortcuts or spurious consensus, and offers a stable reward source without introducing an additional LLM or optimizer. In this way, it effectively overcomes reward hacking and prevents training collapse by implicitly seeking a temporal invariance for true reasoning.
Remark 1. Overall, the two instantiations of Co-rewarding embody our core idea from different perspectives: I leverages data-side analogy-invariance to provide cross supervision, while II employs model-side self-distillation to stabilize learning. Together, they reflect that stable self-supervised reasoning elicitation can emerge from both the diversity of data perspectives and the disentanglement of supervision signals. Moreover, Co-rewarding offers a flexible framework, in which key components, such as pseudo-labeling strategies, data rephrasing techniques, teacher model update rules, and policy optimization, can be seamlessly substituted with other advanced approaches (Yu et al., 2025).
4 Experiments
4.1 Setups
Backbone Models and Baselines. We employ a diverse set of LLMs from different families and scales in our experiments, including the Qwen2.5 series (Qwen2.5-3B/7B) (Qwen et al., 2025), the Qwen3 series (Qwen3-1.7B/4B/8B-Base) (Yang et al., 2025), and the Llama3 series (Llama-3.2-3B-Instruct) (Meta, 2024). Beyond the vanilla GRPO that utilized the GT label for rewarding, we compare our Co-rewarding against several recent state-of-the-art (SoTA) self-reward reasoning approaches, denoted as Self-Certainty (Zhao et al., 2025b), Entropy (Prabhudesai et al., 2025) and Majority Voting (Shafayat et al., 2025). The details of all baselines are summarized in Appendix C.1.
Implementation Details. We implement our algorithms based on the VeRL framework (Sheng et al., 2024), and experiments are conducted on 4 $\times$ H100-80GB GPUs. For our experiments, we totally use three training sets: MATH (Hendrycks et al., 2021) (7,500 questions), DAPO-14k (Yu et al., 2025) (en-version of DAPO-Math-17k, about 14.1k questions), and OpenRS (Dang & Ngo, 2025) (7,000 questions). During RL training, we use a global batch size of 128, set the number of rollouts to $G=\tilde{G}=8$ per question for both I and II, and adopt AdamW with a learning rate of $3\times 10^{-6}$. In Co-rewarding-I, question rephrasing is performed by the open-source Qwen3-32B model. In Co-rewarding-II, the EMA weight is scheduled from $\alpha_{\textrm{start}}=0.99$ to $\alpha_{\textrm{end}}=0.9999$ using cosine annealing. More implementation details are reported in Appendix C.2.
Table 1: Main Results (%) of Co-rewarding and baselines trained on MATH. Cell background colors indicate relative performance: darker colors denote better results within each model group. Additional results of Qwen2.5-3B/7B and Qwen3-1.7B-Base trained on MATH refer to Table 7.
Training Set: MATH Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Before RL 72.4 27.82 20.93 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 82.6 87.26 54.22 30.52 63.25 52.78 57.11
- Self-Certainty (Zhao et al., 2025b) 80.2 80.74 50.75 27.20 64.38 50.98 54.17
- Entropy (Prabhudesai et al., 2025) 80.2 87.19 49.54 29.38 62.00 51.81 54.86
- Majority-Voting (Shafayat et al., 2025) 79.8 89.76 49.09 30.52 63.38 51.80 56.93
- Co-rewarding-I (Ours) 81.2 93.70 51.20 30.81 66.00 55.79 59.95
- Co-rewarding-II (Ours) 80.8 92.42 53.46 30.23 62.83 60.70 57.50
Qwen3-4B-Base
Before RL 71.2 26.15 21.08 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 78.6 89.76 51.20 26.07 55.38 47.80 53.96
- Self-Certainty (Zhao et al., 2025b) 71.6 71.79 38.86 22.37 57.00 48.15 48.93
- Entropy (Prabhudesai et al., 2025) 77.0 88.10 47.44 25.59 52.88 50.44 49.90
- Majority-Voting (Shafayat et al., 2025) 77.4 90.07 45.33 26.54 57.50 48.78 54.35
- Co-rewarding-I (Ours) 78.8 91.28 46.08 26.64 56.50 50.35 53.26
- Co-rewarding-II (Ours) 78.0 88.86 45.93 26.25 55.00 51.30 53.88
Llama-3.2-3B-Instruct
Before RL 39.2 65.73 10.54 9.86 25.37 57.32 31.14
- GT-Reward (Shao et al., 2024) 47.0 77.94 22.14 9.57 31.87 47.51 34.32
- Self-Certainty (Zhao et al., 2025b) 43.4 74.91 18.83 9.95 25.87 54.88 33.34
- Entropy (Prabhudesai et al., 2025) 43.4 66.19 20.18 11.66 24.62 54.70 33.52
- Majority-Voting (Shafayat et al., 2025) 46.8 78.77 20.48 11.00 31.25 47.96 33.18
- Co-rewarding-I (Ours) 50.2 79.45 23.80 11.28 29.88 48.89 33.77
- Co-rewarding-II (Ours) 49.8 79.30 22.59 10.80 30.63 49.90 33.61
Evaluation Details. To provide a comprehensive evaluation of model capabilities, we utilize a diverse set of benchmarks spanning mathematical reasoning, code generation, instruction-following, and general multi-task abilities. Specifically: (1) Mathematical reasoning: MATH500 (Lightman et al., 2024), GSM8K (Cobbe et al., 2021), and AMC (Li et al., 2024a). (2) Code generation: LiveCodeBench (Jain et al., 2025) release_v6 and CRUX (Gu et al., 2024). (3) Instruction-following and multi-task abilities: IFEval (Zhou et al., 2023b) and MMLU-Pro (Wang et al., 2024). Additional evaluation details are provided in Appendix C.3.
4.2 Experimental Results
4.2.1 Main Performance of Co-rewarding
Superior Performance of Co-rewarding over self-rewarding baselines. Table 1 and Table 2 report the experimental results trained on MATH and DAPO-14k, respectively. We observe that both Co-rewarding-I and Co-rewarding-II occupy more darker cells in the tables, demonstrating stronger performance than other self-rewarding SoTA baselines. Specifically, Co-rewarding-I achieves an average relative performance gain of $3.46\%$ over the best baselines across three mathematical benchmarks and models in Table 1, while Co-rewarding-II achieves a larger average relative gain of $7.29\%$ in Table 2. Additional results on other training sets and LLMs are provided in Appendix D.1.
Surpassing GT-Reward in certain benchmarks. Surprisingly, we observe that both Co-rewarding-I and Co-rewarding-II outperform GT-Reward in certain cases. On GSM8K, they together achieve an average relative improvement of $2.94\%$ over GT-Reward in Table 1, while Co-rewarding-II further delivers a larger gain of $5.44\%$ in Table 2. Co-rewarding also shows advantages on the coding benchmark CRUX in several cases. Notably, Co-rewarding-II achieves a remarkably high Pass@1 of 94.01% with Qwen3-8B-Base on GSM8K. These findings highlight the potential of self-supervised RL to elicit stronger reasoning abilities, particularly with Co-rewarding mitigating training collapse.
Table 2: Main Results (%) of Co-rewarding and baselines trained on DAPO-14k. Cell background colors indicate relative performance: darker colors denote better results within each model group. Additional Results of Qwen3-8B-Base trained on OpenRS refer to Table 8.
Training Set: DAPO-14k Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Before RL 72.4 27.82 20.93 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 86.6 87.19 61.75 30.52 63.75 53.11 60.27
- Self-Certainty (Zhao et al., 2025b) 82.0 77.63 49.85 27.77 60.75 50.58 54.24
- Entropy (Prabhudesai et al., 2025) 79.4 80.82 45.48 30.14 62.00 51.56 54.57
- Majority-Voting (Shafayat et al., 2025) 78.6 91.66 50.00 30.33 61.62 51.54 55.65
- Co-rewarding-I (Ours) 78.4 88.02 51.20 29.38 62.50 50.17 55.39
- Co-rewarding-II (Ours) 80.6 94.01 54.37 31.66 67.12 53.31 59.83
Qwen3-4B-Base
Before RL 71.2 26.15 21.08 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 83.6 85.14 52.86 18.58 56.88 47.70 55.35
- Self-Certainty (Zhao et al., 2025b) 68.4 44.81 35.39 25.88 50.12 45.58 48.84
- Entropy (Prabhudesai et al., 2025) 76.6 82.79 43.37 26.35 50.75 48.20 50.22
- Majority-Voting (Shafayat et al., 2025) 73.4 64.06 40.81 26.16 53.00 48.91 51.06
- Co-rewarding-I (Ours) 73.8 75.89 43.83 26.25 50.12 46.84 51.51
- Co-rewarding-II (Ours) 77.8 91.89 48.49 26.64 54.87 48.90 52.83
Llama-3.2-3B-Instruct
Before RL 39.2 65.73 10.54 9.86 25.37 57.32 31.14
- GT-Reward (Shao et al., 2024) 49.4 78.17 25.90 10.33 31.37 53.10 33.83
- Self-Certainty (Zhao et al., 2025b) 42.4 74.71 17.32 11.18 28.38 54.50 33.51
- Entropy (Prabhudesai et al., 2025) 44.0 65.85 17.32 9.95 25.00 55.78 31.95
- Majority-Voting (Shafayat et al., 2025) 42.8 70.96 17.62 10.14 29.50 54.07 32.95
- Co-rewarding-I (Ours) 46.0 70.58 20.93 9.57 27.25 53.04 32.61
- Co-rewarding-II (Ours) 49.8 78.62 19.73 10.43 32.25 51.92 34.46
Figure 3: Performance curves comparison on validation set. Top: Qwen3-1.7B-Base and Qwen2.5-7B trained on the MATH set. Bottom: Qwen3-8B-Base and Llama-3.2-3B-Instruct trained on the DAPO-14k set.
Figure 4: Reward (left) and response length (right) of Qwen3-8B-Base and Llama-3.2-3B-Instruct trained on DAPO-14k. Due to different reward scale from other methods, Entropy reward is plotted on the right $y$-axis of left panels, where the reward is the negative entropy.
Code generalization with preserved general performance. Although trained solely on math-oriented datasets, the models show improvements on coding benchmarks, suggesting a cross-domain generalization from math to code in self-supervised reasoning elicitation. Moreover, Co-rewarding preserves general instruction-following and multi-task ability on MMLU-Pro and IFEval. As shown in Table 4, Co-rewarding-II outperforms other self-rewarding baselines in 12 of 14 MMLU-Pro categories, demonstrating that its gains do not come at the expense of broader general-domain performance. More detailed results of MMLU-Pro and IFEval refer to Appendix D.7 and D.8.
Importance of stability for performance gain. As shown in Table 2, self-rewarding baselines exhibit noticeably limited performance gain in certain cases, such as Self-Certainty with Qwen3-4B-Base on GSM8K. Figure 5 further reflects this by showing that baselines improve quickly at the beginning but soon collapse on GSM8K and AMC, whereas Co-rewarding sustains steady progress. This collapse restricts the baselines to effective training on only a small portion of the data, preventing further improvements with continued training. These observations underscore the importance of avoiding training collapse in self-supervised RL to unlock further performance gains.
Figure 5: Performance and Stability on GSM8K and AMC. The gains of Co-rewarding arise from its training stability, which supports continuous improvements throughout learning.
Table 3: Ablation study of Co-rewarding. For Co-rewarding-I, ablations train only on original or rephrased data. For Co-rewarding-II, ablation removes EMA updates of the reference teacher.
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Co-rewarding-I 81.2 93.70 51.20 30.81 66.00 55.79 59.95
- Majority-Voting w/ Original 79.8 89.76 49.09 30.52 63.38 51.80 56.93
- Majority-Voting w/ Rephrased 79.2 91.51 50.75 31.66 60.38 52.24 57.26
Co-rewarding-II 80.6 94.01 54.37 31.66 67.12 53.31 59.83
- w/o Updating Reference 78.0 88.40 51.66 30.62 63.75 52.48 58.01
Llama-3.2-3B-Instruct
Co-rewarding-I 50.2 79.45 23.80 11.28 29.88 48.89 33.77
- Majority-Voting w/ Original 46.8 78.77 20.48 11.00 31.25 47.96 33.18
- Majority-Voting w/ Rephrased 44.0 78.85 21.23 10.04 17.25 47.84 33.72
Co-rewarding-II 49.8 78.62 19.73 10.43 32.25 51.92 34.46
- w/o Updating Reference 45.0 76.72 17.92 10.05 30.63 51.33 33.94
4.2.2 Further Analysis
Co-rewarding alleviates collapse and provides stable self-supervised RL. We use 5,000 questions from the MATH test split as a validation set to monitor training process. Figure 4 shows the all three self-rewarding baselines collapse on both MATH and DAPO-14k. Co-rewarding-I remains stable on MATH but still collapses on DAPO-14k, suggesting that its stability depends on the property of training data. A plausible explanation is that MATH contains more questions with rich background descriptions than DAPO-14k, which makes rephrasing more effective in amplifying differences between original and rephrased data, thereby strengthening contrastive agreement. In contrast, Co-rewarding-II consistently maintains stability across datasets, as its design decouples supervision from the online policy and thus breaks the entanglement between supervision signals and the policy itself.
Co-rewarding attempts to balance exploration-exploitation. Figure 4 shows reward and response length curves. Entropy and Majority-Voting quickly reach the highest reward, indicating reward hacking rather than genuine reasoning improvement. In contrast, GT-Reward and Co-rewarding exhibit smoother, gradually increasing rewards, reflecting stable training. The response length curves further illustrate this difference: GT-Reward lengthens responses to explore correct reasoning paths; Majority-Voting collapses to short outputs, restricting exploration; and Entropy collapses its probability mass onto a small set of tokens, repeatedly generating them until truncation. Co-rewarding instead maintains moderate response lengths throughout training, suggesting a balanced exploration–exploitation trade-off. Additional curves for other LLMs are provided in Appendix D.2.
Table 4: Detailed performance of MMLU-Pro with Qwen3-8B-Base trained on DAPO-14k. More results refer to Appendix D.7.
MMLU-Pro (Qwen3-8B-Base)
Methods biology business chemistry computer sci. economics health history
- GT-Reward 77.96 70.85 60.42 61.95 71.33 59.79 51.44
- Self-Certainty 75.73 58.05 50.53 56.83 69.31 54.77 50.40
- Entropy 74.76 59.70 51.33 56.10 67.90 55.87 48.04
- Majority-Voting 75.32 61.47 54.24 58.29 69.67 58.20 49.34
- Co-rewarding-I 76.85 61.22 53.45 59.02 66.82 55.62 48.29
- Co-rewarding-II 76.71 68.69 64.58 61.71 68.25 56.85 51.71
Methods law math other philosophy physics psychology engineering
- GT-Reward 31.52 73.28 56.28 52.71 61.97 67.30 46.14
- Self-Certainty 30.43 63.06 51.63 46.29 51.73 66.42 41.07
- Entropy 28.97 63.96 51.51 48.90 53.04 66.80 42.32
- Majority-Voting 31.16 64.62 52.27 48.90 53.27 66.92 40.97
- Co-rewarding-I 30.34 66.17 51.73 48.90 55.19 66.42 39.63
- Co-rewarding-II 31.16 72.17 52.49 52.10 63.21 68.17 47.16
Figure 6: Pseudo label accuracy comparison of Co-rewarding-II, “w/o updating reference”, and Majority-Voting.
Figure 7: Case study: An example comparing the generations from Majority-Voting, Entropy, and our proposed Co-rewarding. The results clearly reveal the reward hacking behavior exhibited by Majority-Voting and Entropy, while ours generate the correct answer. Full results refer to Appendix D.10.
Each part contributes to Co-rewarding. Table 5 summarizes the ablations. For Co-rewarding-I, replacing cross supervision with training only on original or rephrased data yields comparable results, but both underperform the full method. This indicates that the quality of original and rephrased data is similar, while the cross supervision across views provides the crucial gains. For Co-rewarding-II, removing the EMA update of the reference teacher model causes clear degradation, highlighting the necessity of teacher updates for improving pseudo-label quality.
EMA is essential in Co-rewarding-II for improving pseudo-label quality. Figure 6 compares pseudo-label accuracy across Co-rewarding-II, “w/o Updating Reference”, and Majority-Voting. Co-rewarding-II steadily improves accuracy as training progresses, while “w/o Updating Reference” remains nearly flat, underscoring the role of EMA updates in allowing the teacher to co-evolve with the policy and generate higher-quality pseudo labels. By contrast, Majority-Voting briefly improves but then collapses to near zero, evidencing reward hacking through consistent yet incorrect outputs.
Case Study of the model reasoning with different learning methods. Figure 7 provides a concrete example to illustrate the qualitative difference between self-rewarding baselines and our Co-rewarding. Majority-Voting exhibits reward hacking by boxing an incorrect answer “0” to pursue consensus, even though the reasoning steps are correct. Entropy produces repetitive outputs as its decoding probability distribution collapses onto a narrow set of tokens during entropy minimization. In contrast, Co-rewarding generates coherent step-by-step reasoning and correctly boxes the final answer, showing its capacity to overcome reward hacking and elicit genuine reasoning.
5 Conclusion
In this work, we introduced Co-rewarding, a self-supervised RL framework that elicits the reasoning capability of LLMs through complementary supervision. Unlike prior self-rewarding methods that entangle rewards with single-view outputs and risk collapse, Co-rewarding establishes stability by decoupling the reward signal from the current online policy with the single-view output. Specifically, Co-rewarding-I leverages contrastive agreement across semantically analogous questions, while Co-rewarding-II employs a dynamically updated teacher to provide insulated pseudo-labels. Together, these designs construct cross-referable reward signals without explicit labels, aligning RL with invariances in reasoning rather than the mere correctness of isolated outputs. We hope this work will inspire further exploration into self-supervised RL for reasoning to advance the development.
Appendix A Related Work
Large Language Model Reasoning. LLMs have shown impressive performance on vast tasks that require reasoning, including solving mathematical problems, writing code, and answering logical questions. One of the key techniques that has improved LLM reasoning is Chain-of-Thought (CoT) prompting (Wei et al., 2022). CoT encourages the model to generate intermediate reasoning steps before producing the final answer, which has been shown to enhance performance on tasks like arithmetic, commonsense reasoning, and symbolic reasoning. Subsequent work has extended CoT by integrating it with various strategies, including compositional generalization (Zhou et al., 2023a; Khot et al., 2023) and employing structural reasoning approaches (Yao et al., 2023a; Besta et al., 2024; Yang et al., 2024). In addition, CoT serves as a fundamental framework for techniques like fine-tuninig (Zelikman et al., 2022), argentic workflow (Yao et al., 2023b), and paving the way for inference-time scaling (Snell et al., 2024).
RL for Large Language Models. Several RL algorithms have been developed primarily for alignment tasks. Specifically, DPO (Rafailov et al., 2023), CPO (Xu et al., 2024), and their variants (Li et al., 2024b; Guo et al., 2024; Munos et al., 2024; Hong et al., 2024; Xie et al., 2024) rely on pairs of outputs labeled by human preference. In contrast, KTO (Ethayarajh et al., 2024) and BCO (Jung et al., 2024) require only a single binary label (like or dislike) for each output. Besides, the PRM (Uesato et al., 2022; Lightman et al., 2024) and Step-KTO (Lin et al., 2025a) offer step-by-step guidance by incorporating feedback at each reasoning step rather than focusing solely on the final outputs. Recently, the follow-up work of GRPO improves the optimization objective, e.g., DAPO (Yu et al., 2025), Dr. GRPO (Liu et al., 2025a), REINFORCE++ (Hu, 2025), CPPO (Lin et al., 2025b), and GPG (Chu et al., 2025). Another line of research generalizes GRPO to broader applications such as multimodal reasoning (Zhou et al., 2025; Huang et al., 2025; Chu et al., 2025; Liu et al., 2025b; Zhang et al., 2025a) and logical reasoning (Xie et al., 2025).
RL without External Reward. RL methods have shown promising scaling capabilities to enhance the reasoning abilities of LLMs (Guo et al., 2025), yet they are often limited by the availability of training data for reward signals (Gao et al., 2023; Liu et al., 2023). Notably, Wang et al. (Wang et al., 2025) demonstrate that RL can effectively bootstrap LLM reasoning with as little as a single training example, highlighting the potential to minimize or even eliminate reliance on external reward signals during training. Recent efforts leverage distinct strategies for reward assignment. For instance, SIRLC (Pang et al., 2024) and AZR (Zhao et al., 2025a) utilize an LLM-as-the-judge approach to assign rewards. In contrast, methods like SRT, TTRL, and their variants (Shafayat et al., 2025; Zuo et al., 2025; Fang et al., 2025; Zhang et al., 2025b) employ self-consistency (Wang et al., 2022) to generate pseudo-rewards, reducing dependence on external annotations. Meanwhile, INTUITOR, RLSC, and RENT (Zhao et al., 2025b; Li et al., 2025; Prabhudesai et al., 2025) harness the internal confidence scores of LLMs as intrinsic reward signals. Additionally, EMPO and its variants (Zhang et al., 2025c; Agarwal et al., 2025) promote reasoning by minimizing entropy during the reasoning process, further diversifying the approaches to incentivize robust LLM performance.
Algorithm 1 Co-rewarding-I
1:Input: policy model $\pi_{\theta}$, learning rate $\eta$, training dataset $\mathcal{D}$, rephrased training dataset $\mathcal{D^{\prime}}$, total iterations $K$.
2:Output: trained policy model $\pi_{\theta}$.
3:for all iteration $k=1,\dots,K$ do
4: Sample mini-batch inputs $\mathcal{B}\subseteq\mathcal{D}$ and $\mathcal{B}^{\prime}\subseteq\mathcal{D}^{\prime}$.
5:for all input question $x\in\mathcal{B}$ and $x^{\prime}\in\mathcal{B}^{\prime}$ do
6:  Sample rollouts $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\textrm{old}}}(\cdot\mid x)$.
7:  Sample rollouts $\{y^{\prime}_{i}\}_{i=1}^{G^{\prime}}\sim\pi_{\theta_{\textrm{old}}}(\cdot\mid x^{\prime})$.
8:  Obtain pseudo labels by Eq. (8).
9:  Estimate relative advantages by Eq. (7).
10:  Compute the objective by Eq. (6).
11:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{J}_{\mathrm{Co\text{-}rewarding\text{-}I}}(\theta)$.
12:end for
13:end for
Algorithm 2 Co-rewarding-II
1:Input: policy model $\pi_{\theta}$, learning rate $\eta$, training dataset $\mathcal{D}$, total iterations $K$.
2:Output: trained policy model $\pi_{\theta}$.
3:for iteration $k=1,\dots,K$ do
4: Sample mini-batch $\mathcal{B}\subseteq\mathcal{D}$.
5:for all $x\in\mathcal{B}$ do
6:  Sample rollouts $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}^{(k)}(\cdot\mid x)$.
7:  Update the reference teacher by Eq. (11).
8:  Sample rollouts $\{\tilde{y}_{j}\}_{j=1}^{\tilde{G}}\sim\tilde{\pi}_{\mathrm{ref}}^{(k)}(\cdot\mid x)$.
9:  Obtain pseudo label from $\{\tilde{y}_{j}\}_{j=1}^{\tilde{G}}$ by Eq. (10).
10:  Estimate the relative advantage by Eq. (10).
11:  Compute the objective by Eq. (9).
12:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{J}_{\mathrm{Co\text{-}rewarding\text{-}II}}^{(k)}(\theta)$.
13:end for
14:end for
Appendix B Pseudo Code of Co-rewarding
To intuitively present the pipeline of Co-rewarding, we summarize the pseudo codes of Co-rewarding-I and Co-rewarding-II in Algorithm 1 and Algorithm 2, respectively.
Appendix C Additional Experimental Details
C.1 Details of Baselines
We compare our proposed Co-rewarding-I and II against GT-reward and several recent state-of-the-art (SoTA) self-reward approaches:
  • GT-Reward (Shao et al., 2024): Originally introduced by DeepSeek-R1 (Guo et al., 2025), GT-Reward supervises training using ground-truth (GT) answers, determining whether model rollouts are correct or not, to guide RL optimization.
  • Self-Certainty (Zhao et al., 2025b): This method maximizes self-certainty, defined as the KL-divergence between the uniform distribution and the model’s decoding distribution, serving as reward to encourage more confident predictions.
  • Entropy (Prabhudesai et al., 2025): This method minimizes the entropy of the model’s rollout distribution, using negative entropy as reward to maximize model confidence.
  • Majority-Voting (Shafayat et al., 2025): By generating multiple rollouts per question, Majority-Voting selects the most frequent answer as a pseudo-label to supervise training.
For all methods, we adopt the widely used GRPO as the policy optimization algorithm.
C.2 More Implementation Details
The detailed training configurations are summarized in Table 5, and all baseline methods are trained under the same setup for fairness. For the training system prompt, we adopt the official default prompt provided by VeRL [^1] https://github.com/volcengine/verl, shown below:
[…]
In addition, the semantically analogical questions used in Co-rewarding-I are generated by Qwen3-32B through a rewriting prompt. The exact rewriting instruction is provided as follows:
You are given a math problem. Please rewrite it using different wording and a different real-world scenario, while keeping the underlying mathematical meaning and answer exactly the same.
Guidelines:
1. Do not change the math logic or the final answer.
2. Use different words and a new context to make it look like a different problem.
3. Avoid copying phrases or sentence structures from the original.
4. Make sure the rewritten question is natural, clear, and solvable.
5. Output ONLY between the following markers, and strictly in this format (no extra explanation):
### RESULT_START
ORIGINAL:
<original question>
REWRITE:
<rewritten question>
### RESULT_END
Table 5: Detailed training settings.
Settings Co-rewarding-I Co-rewarding-II
Batch Size 128 128
Max Prompt Length 512 512
Max Response Length 3072 3072
Train Steps 170-220 300-330
Learning Rate 3e-6 3e-6
# Policy Rollout $G$ 8 8
# Reference Rollout $\tilde{G} -$ 8
Clip Ratio 0.2 0.2
Warmup Style Cosine Cosine
Warmup Steps Ratio 0.1 0.1
KL Loss Coefficient 0.005 0.001
Optimizer AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$)
Training Temperature 1.0 1.0
Evaluation Temperature 0.8 0.8
EMA $\alpha_{\textrm{start}}$ - 0.99
EMA $\alpha_{\textrm{end}}$ - 0.9999
Table 6: Statistics and usages of datasets used in our experiments.
Dataset Name # Data Size Usage
MATH-Train (Hendrycks et al., 2021) 7,500 Training Set
MATH-Test (Hendrycks et al., 2021) 5,000 Validation Set
DAPO-14k (Yu et al., 2025) 14,109 Training Set
Open-RS (Dang & Ngo, 2025) 7,000 Training Set
MATH500 (Lightman et al., 2024) 500 Evaluation Benchmark
GSM8K (Cobbe et al., 2021) 1,319 Evaluation Benchmark
AMC (Li et al., 2024a) 83 Evaluation Benchmark
LiveCodeBench (Jain et al., 2025) 1,055 Evaluation Benchmark
CRUX (Gu et al., 2024) 800 Evaluation Benchmark
MMLU-Pro (Wang et al., 2024) 12,032 Evaluation Benchmark
IFEval (Zhou et al., 2023b) 541 Evaluation Benchmark
C.3 More Evaluation Details
We conduct the evaluation across a diverse set of benchmarks, spanning mathematical reasoning, code generation, instruction-following, and general multi-task abilities. Specifically: (1) Mathematical reasoning: We evaluate on MATH500 (Lightman et al., 2024), GSM8K (Cobbe et al., 2021), and AMC (Li et al., 2024a). For MATH500 and GSM8K, we report pass@1 accuracy using the lighteval library [^2] https://github.com/huggingface/lighteval. For AMC, we use the ttrl [^3] https://github.com/ruixin31/Spurious_Rewards/tree/main/code/ttrl library and report avg@8 as the metric. (2) Code generation: We assess coding ability using LiveCodeBench (Jain et al., 2025) release_v6 and CRUX (Gu et al., 2024). LiveCodeBench is evaluated with its official evaluation library [^4] https://github.com/LiveCodeBench/LiveCodeBench, and CRUX is evaluated via the ZeroEval library [^5] https://github.com/WildEval/ZeroEval; for both datasets, we report pass@1 accuracy. (3) Instruction-following and multi-task abilities: We evaluate on IFEval (Zhou et al., 2023b) and MMLU-Pro (Wang et al., 2024), using the lm-evaluation-harness library [^6] https://github.com/EleutherAI/lm-evaluation-harness for both. Overall, we summarize the statistics of the datasets used in this paper in Table 6.
Table 7: Supplement Results (%) of Co-rewarding and baselines trained on MATH. Cell background colors: darker colors denote better results within each model group.
Training Set: MATH Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen2.5-3B
Before RL 53.6 19.48 10.69 9.95 18.50 29.83 32.50
- GT-Reward (Shao et al., 2024) 65.4 82.18 32.98 13.93 32.12 33.66 36.74
- Self-Certainty (Zhao et al., 2025b) 64.2 80.52 28.92 10.90 29.00 32.22 33.88
- Entropy (Prabhudesai et al., 2025) 63.2 80.44 29.67 9.05 29.00 32.94 35.35
- Majority-Voting (Shafayat et al., 2025) 64.6 82.41 33.13 14.03 36.38 35.19 35.50
- Co-rewarding-I (Ours) 65.4 84.53 30.57 16.40 36.88 33.86 36.38
- Co-rewarding-II (Ours) 65.2 81.72 32.38 22.25 40.25 32.74 30.79
Qwen2.5-7B
Before RL 69.4 24.71 15.81 3.79 26.38 38.19 44.76
- GT-Reward (Shao et al., 2024) 76.4 88.02 45.63 15.92 45.12 41.49 41.12
- Self-Certainty (Zhao et al., 2025b) 72.8 84.31 38.55 12.04 54.12 37.24 43.30
- Entropy (Prabhudesai et al., 2025) 72.2 81.43 39.61 16.49 51.88 40.33 42.79
- Majority-Voting (Shafayat et al., 2025) 74.4 84.53 40.96 15.45 51.00 38.60 43.35
- Co-rewarding-I (Ours) 74.6 89.61 41.27 15.73 55.58 42.86 40.51
- Co-rewarding-II (Ours) 73.6 89.31 42.77 8.25 47.50 41.82 37.45
Qwen3-1.7B-Base
Before RL 57.0 19.56 8.43 4.45 7.50 33.65 33.00
- GT-Reward (Shao et al., 2024) 69.6 81.57 35.54 13.74 35.25 36.16 39.12
- Self-Certainty (Zhao et al., 2025b) 58.2 40.25 23.04 9.86 18.00 32.96 35.13
- Entropy (Prabhudesai et al., 2025) 63.6 71.79 31.63 13.74 31.37 35.37 36.67
- Majority-Voting (Shafayat et al., 2025) 65.2 81.57 34.78 13.08 34.25 35.45 36.00
- Co-rewarding-I (Ours) 67.6 83.01 32.22 13.50 32.38 35.56 35.53
- Co-rewarding-II (Ours) 66.2 80.89 33.28 14.40 32.88 36.94 37.59
Table 8: Supplement Results (%) of Co-rewarding and baselines trained on OpenRS. Cell background colors: darker colors denote better results within each model group.
Training Set: Open-RS Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Before RL 72.40 27.82 20.93 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 80.20 89.76 54.97 39.00 63.00 52.94 55.49
- Self-Certainty (Zhao et al., 2025b) 82.60 85.22 50.00 37.00 64.62 52.12 56.03
- Entropy (Prabhudesai et al., 2025) 80.60 87.41 48.95 38.00 61.25 52.53 56.80
- Majority-Voting (Shafayat et al., 2025) 78.00 84.23 51.96 36.75 58.00 51.13 54.92
- Co-rewarding-I (Ours) 78.20 92.65 50.60 28.91 63.12 53.11 57.21
- Co-rewarding-II (Ours) 80.00 90.90 53.01 39.75 62.75 52.92 56.55
Qwen3-4B-Base
Before RL 71.20 26.15 21.08 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 78.80 85.22 49.55 33.50 55.12 46.41 50.12
- Self-Certainty (Zhao et al., 2025b) 73.20 33.43 35.84 32.50 49.50 46.47 48.24
- Entropy (Prabhudesai et al., 2025) 76.80 87.57 42.62 35.00 53.87 47.61 52.42
- Majority-Voting (Shafayat et al., 2025) 76.00 64.14 44.58 32.25 50.25 46.35 48.75
- Co-rewarding-I (Ours) 72.80 83.93 39.41 26.54 53.25 48.11 50.82
- Co-rewarding-II (Ours) 76.60 89.23 42.32 34.00 51.50 48.45 51.80
Figure 8: Performance curves on validation set. Left to Right: {Qwen2.5-3B, Qwen3-4B-Base, Qwen3-8B-Base, Llama-3.2-3B-Instruct} trained on MATH, Qwen3-4B-Base trained on DAPO-14k.
Appendix D Additional Experimental Results
D.1 More Results on Other Training Sets and LLMs
Table 7 reports additional results of Qwen2.5-3B and Qwen3-1.7B-Base and Qwen3-4B-Base trained on MATH, while Table 8 extends the experiments of Qwen3-8B-Base and Qwen3-4B-Base to another training set OpenRS (Dang & Ngo, 2025). It can be observed that Co-rewarding occupies relatively darker areas. Across models and training sets, Co-rewarding-I and II achieve an average relative improvement of $2.23\%$ on GSM8K, with notably high pass@1 scores of $92.65\%$ and $90.90\%$ for Qwen3-8B-Base trained on OpenRS, respectively. Moreover, thanks to its stability, Co-rewarding-II delivers more reliable gains than self-rewarding baselines, which occasionally suffer lower performance on certain models or benchmarks, e.g., Self-Certainty on Qwen3-1.7B-Base in Table 7 or Majority-Voting on Qwen3-4B-Base in Table 8. These results further demonstrate the effectiveness of Co-rewarding.
Table 9: Performance (%) of test-time training (TTT). Since self-supervised methods are label-free, they can be leveraged during inference for test-time training to further enhance performance.
LLMs Methods AMC
avg@8 pass@8 avg@16 pass@16 avg@32 pass@32 avg@64 pass@64
Qwen2.5-7B Before-TTT 15.81 46.99 17.55 66.27 16.34 74.70 17.32 75.90
Self-Certainty 41.57 74.70 39.23 74.70 39.68 78.31 39.95 87.95
Entropy 38.70 56.63 39.76 68.67 39.57 79.52 39.34 81.93
Majority-Voting 43.67 63.86 43.67 67.47 43.49 78.31 44.35 85.54
Co-rewarding-I 44.88 60.24 45.33 60.24 45.44 71.08 45.76 73.49
Co-rewarding-II 43.22 69.88 41.34 75.90 40.36 78.31 41.64 87.95
Qwen3-8B-Base Before-TTT 20.93 61.45 21.31 73.49 19.58 79.52 20.97 86.75
Self-Certainty 49.85 78.31 50.68 78.31 50.41 84.34 49.55 89.16
Entropy 48.64 74.70 49.92 80.72 49.96 87.95 50.23 89.16
Majority-Voting 50.90 73.49 50.00 72.29 50.60 80.72 51.36 85.54
Co-rewarding-I 52.86 68.67 53.46 74.70 53.24 81.93 53.58 84.34
Co-rewarding-II 48.64 72.29 48.19 73.49 50.19 83.13 49.28 91.57
Figure 9: Curves of reward (Left), response length (Middle), and pseudo label accuracy (Right) of Qwen3-4B-Base trained on DAPO-14k. Entropy reward is plotted on the right $y$-axis due to its different reward scale. Note that entropy minimization is to maximizing the negative entropy.
Table 10: Detailed MMLU-Pro performance on Qwen3-4B-Base and Llama-3.2-3B-Instruct traind on DAPO-14k. Results are reported for each of the 14 categories in MMLU-Pro.
Qwen3-4B-Base
Methods biology business chemistry computer sci. economics engineering health
- GT-Reward 73.50 63.49 59.71 56.34 65.05 42.93 50.86
- Self-Certainty 71.41 54.37 45.93 50.73 63.27 35.91 50.12
- Entropy 70.99 56.02 50.44 48.29 63.15 34.37 48.41
- Majority-Voting 70.43 55.77 52.83 53.41 62.79 38.09 50.61
- Co-rewarding-I 73.92 59.82 50.71 54.15 64.93 41.49 49.76
- Co-rewarding-II 72.66 59.95 55.65 53.41 64.10 39.73 50.61
Methods history law math other philosophy physics psychology
- GT-Reward 44.88 26.34 69.80 48.81 44.69 57.04 65.79
- Self-Certainty 39.63 24.43 59.44 43.94 40.08 47.04 59.65
- Entropy 40.68 26.43 60.99 45.13 43.69 50.89 61.90
- Majority-Voting 40.94 23.43 64.17 43.39 44.09 50.73 63.66
- Co-rewarding-I 40.94 23.25 63.73 44.91 42.69 50.58 60.78
- Co-rewarding-II 42.26 24.79 67.58 44.59 41.88 54.19 62.91
Llama3.2-3B-Instruct
Methods biology business chemistry computer sci. economics engineering health
- GT-Reward 54.81 36.25 25.18 33.41 42.65 21.57 39.36
- Self-Certainty 55.23 32.95 27.21 31.95 42.77 20.54 39.12
- Entropy 52.86 31.05 23.94 32.93 41.71 20.43 38.02
- Majority-Voting 56.07 32.95 22.79 30.98 44.19 18.99 39.61
- Co-rewarding-I 51.88 34.22 22.88 34.88 44.67 19.09 38.63
- Co-rewarding-II 56.21 34.35 27.03 35.61 43.01 19.92 40.34
Methods history law math other philosophy physics psychology
- GT-Reward 30.18 22.71 34.20 34.74 32.06 28.33 50.38
- Self-Certainty 30.45 24.98 33.38 31.60 29.86 28.56 50.50
- Entropy 33.86 21.89 32.35 33.01 29.46 24.25 47.50
- Majority-Voting 32.02 25.25 34.35 34.20 29.86 24.79 48.25
- Co-rewarding-I 33.86 23.25 32.12 33.01 31.86 25.40 48.75
- Co-rewarding-II 32.28 24.34 35.83 36.26 33.27 28.18 49.12
D.2 More Curves of Reward, Response Length and Pseuo Label Accuracy
Figure 9 supplements the reward and response curves of Qwen3-4B-Base trained on DAPO-14k. The trends are consistent with Qwen3-8B-Base and Llama-3.2-3B-Instruct in Figure 4: Majority-Voting and Entropy rapidly increase rewards at early stage and quickly peak, a clear sign of reward hacking. In contrast, GT-Reward and Co-rewarding-II exhibit smoother, steadily rising rewards, indicating genuine learning of reasoning ability. Moreover, Co-rewarding-II maintains moderate response lengths on Qwen3-4B-Base, further demonstrating its generality in balancing the exploration–exploitation trade-off during reasoning training, which is a core principle of RL (Wang et al., 2018).
Additionally, the right panel of Figure 9 presents the pseudo-label accuracy of Qwen3-4B-Base, showing trends consistent with Qwen3-8B-Base and Llama-3.2-3B-Instruct in Figure 6. As training progresses, Co-rewarding-II steadily improves pseudo-label accuracy, while “w/o Updating Reference” remains around 25%. Majority-Voting briefly increases accuracy but soon collapses to zero, clearly indicating reward hacking. This highlights our design philosophy of pairing a fast policy student with a slowly updated teacher, which decouples supervision from the online policy while enabling the teacher to co-evolve with the student, thereby sustaining improvements in pseudo-label quality.
D.3 More Results of Validation Performance Curves
As a supplement to Figure 4, Figure 8 presents validation performance curves for Qwen2.5-3B, Qwen3-4B-Base, Qwen3-8B-Base, Llama-3.2-3B-Instruct trained on MATH, as well as Qwen3-4B-Base trained on DAPO-14k. Self-Certainty and Entropy collapse rapidly across all settings, as their supervision signals are tied to internal confidence or entropy and are easily exploited. Majority-Voting also collapses in several cases, reflecting that sampling pseudo labels from outputs cannot prevent hacking. By contrast, Co-rewarding-I maintains stability across MATH-trained models through data-side contrastive agreement, while Co-rewarding-II consistently provides stability across all models and datasets by disentangling supervision with a slowly updated teacher, making hacking substantially harder and optimization more reliable.
D.4 Results of Test-time Training (TTT)
Thanks to the label-free nature of self-supervised methods, which do not require GT labels, they are naturally compatible with test-time training (TTT), enabling further refinement of the model during inference. Table 9 reports the TTT results on the challenging competition-level benchmark AMC across Co-rewarding and other self-rewarding baselines. We observe that Co-rewarding matches or even surpasses existing methods, achieving the best results on 11 out of 18 metrics. These findings broaden the applicability of self-supervised RL: beyond post-training for reasoning elicitation, it can also be leveraged at inference time to further improve performance on specific benchmarks.
Figure 10: Performance curves on benchmarks of MATH500, GSM8K, AMC and CRUX across Qwen3-8B-Base, Qwen3-4B-Base, and Llama-3.2-3B-Instruct trained on DAPO-14k.
D.5 More Results of Benchmark Performance Curves
As a supplement to Figure 4 and Figure 5, Figure 10 presents performance curves on MATH500, GSM8K, AMC, and CRUX with Qwen3-8B-Base, Qwen3-4B-Base, and Llama-3.2-3B-Instruct. Consistent with earlier findings, Self-Certainty, Entropy, and Majority-Voting rapidly collapse across benchmarks and models, while Co-rewarding-II and GT-Reward sustain continued and stable improvements. These results underscore the link between performance and training stability: stable training enables models to continue improving by effectively learning knowledge from more data.
Figure 11: Curves of voting accuracy of Majority-Voting, Co-rewarding-I and its ablations with Qwen2.5-3B, Qwen3-1.7B-Base and Qwen3-8B-Base trained on MATH.
Figure 12: Voting accuracy of Majority-Voting, Co-rewarding-I and its ablated variants across different difficulty levels of questions Top: Qwen3-1.7B-Base. Bottom: Qwen2.5-3B.
D.6 Voting Accuracy Analysis of Co-rewarding-I
To demonstrate the stability and efficiency of Co-rewarding-I, we compare its voting accuracy against that of Majority-Voting in Figure 11 and Figure 12. These experiments are conducted on Qwen2.5-3B, Qwen3-1.7B-Base and Qwen3-8B-Base models, all trained on the MATH dataset. Across all settings, the Majority-Voting method exhibits reward hacking, where its performance sharply declines after reaching an early peak, particularly on more difficult questions (levels 2 to 5). In contrast, Co-Rewarding-I maintains a stable voting accuracy on both original and rephrased data. Ultimately, it achieves the highest overall voting accuracy across all models and dataset configurations.
Table 11: Detailed MMLU-Pro performance on Qwen3-8B-Base and Qwen3-4B-Base trained on OpenRS. Results are reported for each of 14 categories in MMLU-Pro.
Qwen3-4B-Base
Methods biology business chemistry computer sci. economics engineering health
- GT-Reward 70.99 59.82 52.30 54.63 65.05 39.01 51.22
- Self-Certainty 69.87 54.50 44.08 49.27 63.63 37.36 50.24
- Entropy 70.71 58.68 49.03 51.22 63.39 37.46 49.63
- Majority-Voting 69.60 55.77 47.17 53.17 63.39 36.02 48.78
- Co-rewarding-I 69.04 55.39 47.79 53.41 63.86 38.39 50.61
- Co-rewarding-II 70.85 58.81 53.27 53.90 66.11 37.15 52.81
Methods history law math other philosophy physics psychology
- GT-Reward 39.63 24.98 65.58 47.84 40.68 54.50 62.53
- Self-Certainty 39.63 24.25 58.11 46.65 40.88 46.42 61.40
- Entropy 39.90 22.16 62.18 45.02 43.09 50.19 59.90
- Majority-Voting 40.68 22.52 60.25 46.10 41.08 48.42 60.65
- Co-rewarding-I 40.68 24.25 62.18 44.37 44.49 49.58 61.65
- Co-rewarding-II 41.21 25.89 64.91 45.24 39.28 52.27 59.40
Qwen3-8B-Base
Methods biology business chemistry computer sci. economics engineering health
- GT-Reward 74.76 63.24 55.48 63.17 68.96 41.38 57.09
- Self-Certainty 75.03 63.62 53.62 55.61 68.96 39.83 57.09
- Entropy 75.73 64.39 54.51 58.29 65.05 41.69 55.87
- Majority-Voting 76.15 60.20 54.15 56.34 69.91 38.91 55.75
- Co-rewarding-I 76.43 65.78 57.07 62.20 69.43 43.14 56.60
- Co-rewarding-II 76.84 64.25 54.68 62.43 68.12 42.00 58.06
Methods history law math other philosophy physics psychology
- GT-Reward 50.92 30.25 67.58 52.49 51.10 57.20 67.67
- Self-Certainty 49.34 28.88 68.02 51.62 52.10 56.89 66.42
- Entropy 50.39 30.43 65.28 51.41 47.09 54.50 66.67
- Majority-Voting 48.03 28.88 63.43 53.68 48.10 52.50 64.66
- Co-rewarding-I 50.13 29.97 68.54 52.92 50.70 56.66 65.54
- Co-rewarding-II 51.44 30.06 65.80 51.51 52.10 57.58 65.78
D.7 More Results of MMLU-Pro Evaluation
As a complement to Table 4, Table 10 and Table 11 report detailed MMLU-Pro results for models trained on DAPO-14k and OpenRS, respectively. We observe that Co-rewarding consistently preserves general-domain performance across diverse subjects, indicating that though trained on math-oriented datasets, its improvements do not come at the cost of broader capabilities from other domains.
D.8 More Results of IFEval Evaluation
The aim of IFEval is used to evaluate the instruction-following ability of LLMs. In Table 1, Table 2, Table 7 and Table 8, we report average IFEval performance due to space constraints. Specifically, the evaluation of IFEval includes four metrics: {prompt_level_strict_acc, inst_level_strict_acc, prompt_level_loose_acc and inst_level_loose_acc}, which apply different levels of answer matching. As a supplement, complete results are provided in Table 12, Table 13, and Table 14. The results show that Co-rewarding not only preserves the inherent instruction-following ability of base models but also often surpasses GT-Reward across multiple models. This further confirms that Co-rewarding’s gains on mathematical and coding benchmarks are achieved without sacrificing general-domain instruction-following ability.
Table 12: Detailed IFEval Performance on Qwen2.5-3B/7B, Qwen3-1.7B/4B/8B-Base and Llama-3.2-3B-Instruct traind on MATH. Results are reported for loose and strick settings respectively.
Methods IFEval
Average Prompt Strict Prompt Loose Inst. Strict Inst. Loose Average Prompt Strict Prompt Loose Inst. Strict Inst. Loose
Qwen2.5-3B Qwen2.5-7B
Before RL 29.83 22.55 27.17 31.89 37.70 38.19 29.57 34.57 41.85 46.76
- GT-Reward 33.66 25.51 31.42 35.85 41.85 41.49 31.79 39.56 43.65 50.96
- Self-Certainty 32.22 24.40 29.76 34.65 40.05 37.24 28.47 34.38 40.05 46.04
- Entropy 32.94 24.77 30.50 35.13 41.37 40.33 30.13 37.87 43.29 50.00
- Majority-Voting 35.19 26.25 32.72 37.53 44.24 38.60 29.21 35.86 41.61 47.72
- Co-rewarding-I 33.86 23.84 31.61 36.09 43.88 41.73 32.35 39.37 44.48 50.72
- Co-rewarding-II 32.74 23.29 29.02 36.33 42.33 41.82 31.79 40.29 43.88 51.31
Qwen3-1.7B-Base Qwen3-4B-Base
Before RL 33.65 25.69 30.86 36.45 41.60 46.43 36.04 44.18 48.68 56.83
- GT-Reward 36.16 27.35 31.79 40.64 44.84 47.80 37.34 46.77 49.40 57.67
- Self-Certainty 32.96 24.58 29.20 36.69 41.36 48.15 39.37 46.76 49.52 56.95
- Entropy 35.37 26.61 31.42 39.44 44.00 50.44 40.67 48.61 52.52 59.07
- Majority-Voting 35.45 26.06 32.16 38.72 48.84 48.78 37.89 47.50 50.36 59.65
- Co-rewarding-I 35.56 27.91 31.23 39.32 43.76 50.35 40.67 49.35 51.56 59.83
- Co-rewarding-II 36.94 27.17 33.64 40.05 46.88 51.30 41.40 49.54 53.12 61.15
Qwen3-8B-Base Llama3-2-Instruct
Before RL 50.32 40.11 50.27 51.07 59.83 57.32 46.77 55.27 60.19 67.03
- GT-Reward 52.78 41.96 51.76 54.44 62.95 47.41 37.34 42.88 52.52 57.31
- Self-Certainty 50.98 39.74 49.54 52.88 61.75 54.88 43.81 52.68 58.15 64.87
- Entropy 51.81 40.67 51.20 52.76 62.59 54.70 43.81 52.68 57.67 64.63
- Majority-Voting 51.80 39.74 51.02 53.60 62.83 47.96 37.34 43.44 52.88 58.18
- Co-rewarding-I 55.79 43.99 57.11 55.63 66.42 49.14 39.37 45.66 53.12 58.39
- Co-rewarding-II 60.70 55.64 65.59 56.00 65.59 49.90 39.93 45.66 54.68 59.35
Table 13: Detailed IFEval performance on Qwen3-4B/8B-Base and Llama-3.2-3B-Instruct traind on DAPO-14k. Results are reported for loose and strict settings in IFEval, respectively.
Methods IFEval
Average Prompt Strict Prompt Loose Inst. Strict Inst. Loose
Qwen3-4B-Base
Before RL 46.43 36.04 44.18 48.68 56.83
- GT-Reward 47.70 37.52 45.84 49.76 57.67
- Self-Certainty 45.58 35.67 43.99 47.84 54.80
- Entropy 48.20 37.71 46.58 50.48 58.03
- Majority-Voting 48.91 39.19 47.69 50.24 58.51
- Co-rewarding-I 46.84 36.41 45.66 48.80 56.47
- Co-rewarding-II 48.90 39.56 46.21 51.44 58.39
Qwen3-8B-Base
Before RL 50.32 40.11 50.27 51.07 59.83
- GT-Reward 53.11 41.59 52.13 54.56 64.15
- Self-Certainty 50.58 41.04 49.54 51.68 60.07
- Entropy 51.56 41.59 49.91 53.48 61.27
- Majority-Voting 51.54 41.22 51.02 52.64 61.27
- Co-rewarding-I 50.17 40.67 48.24 52.16 59.59
- Co-rewarding-II 53.31 41.40 53.23 54.20 64.39
Llama3.2-3B-Instruct
Before RL 57.32 46.77 55.27 60.19 67.03
- GT-Reward 53.10 42.33 49.91 57.19 62.95
- Self-Certainty 54.50 44.55 51.76 58.03 63.67
- Entropy 55.78 45.29 53.23 59.11 65.47
- Majority-Voting 54.07 42.33 52.50 56.83 64.63
- Co-rewarding-I 53.04 42.33 51.02 55.76 63.07
- Co-rewarding-II 51.92 41.59 48.24 56.00 61.87
Table 14: Detailed IFEval Performance on Qwen3-8B/4B-Base trained on Open-RS. Results are reported for loose and strict settings in IFEval, respectively.
Methods IFEval
Average Prompt Strict Prompt Loose Inst. Strict Inst. Loose Average Prompt Strict Prompt Loose Inst. Strict Inst. Loose
Qwen3-8B-Base Qwen3-4B-Base
Before RL 50.32 40.11 50.27 51.07 59.83 46.43 36.04 44.18 48.68 56.83
- GT-Reward 52.53 41.59 51.02 54.56 62.95 47.80 37.34 46.77 49.40 57.67
- Self-Certainty 52.12 41.59 50.83 53.72 62.35 46.47 35.86 44.73 48.56 56.71
- Entropy 52.94 43.25 51.94 53.72 62.83 46.41 36.97 44.73 48.68 55.28
- Majority-Voting 51.13 40.67 49.35 53.36 61.15 46.35 36.41 44.18 48.80 56.00
- Co-rewarding-I 53.11 41.40 53.05 53.95 64.02 46.83 35.86 46.77 47.84 56.83
- Co-rewarding-II 52.92 42.14 52.50 54.08 62.95 48.45 38.26 46.77 50.24 58.51
Table 15: Case studies: Original vs. Rephrased Questions
Original Question Rephrased Question
Sam is hired for a 20-day period. On days that he works, he earns \$60. For each day that he does not work, \$30 is subtracted from his earnings. At the end of the 20-day period, he received \$660. How many days did he not work? A contractor agrees to a job lasting 20 days. For every day the job is completed on time, the contractor earns \$60. However, for each day the work is delayed, a fine of \$30 is applied. After the 20-day period, the contractor’s total earnings are \$660. How many days was the job delayed?
Karen drove continuously from 9:40 a.m. until 1:20 p.m. of the same day and covered a distance of 165 miles. What was her average speed in miles per hour? A traveler set off at 9:40 a.m. and reached their destination at 1:20 p.m. the same day after traveling a total of 165 miles. What was their average speed during the trip in miles per hour?
Solve for $x$: $\tfrac{1}{2}+\tfrac{1}{x}=\tfrac{5}{6}$. A tank is partially filled by two different pipes. One pipe fills half the tank in an hour, and together with another pipe, they fill five-sixths of the tank in the same time. If the second pipe alone fills $\tfrac{1}{x}$ of the tank in an hour, find the value of $x$.
D.9 Original Questions vs. Rephrased Questions
To provide an intuitive illustration, we present several examples of original questions with their rephrased versions in Table 15. We observe that such rephrasings are reasonable and effective, as they preserve the same underlying mathematical essence while presenting the problems in a substantially different surface form. This reflects the high quality of our rephrased data and forms the basis of Co-rewarding-I: by leveraging contrastive agreement across data-invariant variants, the model is encouraged to elicit more robust reasoning ability.
D.10 Complete Case Study
As a supplement to Figure 7, we present the complete generation outputs of this case study. The full outputs clearly reveal the reward hacking behaviors of existing self-rewarding baselines. Self-Certainty and Entropy fall into repetitive outputs—for example, Self-Certainty repeatedly generates “Understanding,” and Entropy repeatedly produces “Simplify the next fraction” until truncated at the maximum length. This arises because their decoding probability mass collapses onto a small subset of tokens, leading the model to loop over them. Majority-Voting shows another form of reward hacking by boxing an incorrect answer “0” to maximize consensus across rollouts and thereby secure the highest reward. In contrast, both Co-rewarding and GT-Reward yield the correct final answer along with a coherent step-by-step reasoning path.
Question: $\frac{m}{n}$ is the Irreducible fraction value of $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$ , what is the value of $m+n$?
Co-rewarding-I Answer: To solve the problem, we need to simplify the given expression step by step and then find the irreducible fraction form. The expression is: $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}.$ First, we simplify the innermost fraction: $3+\frac{1}{3}=\frac{9}{3}+\frac{1}{3}=\frac{10}{3}.$ Now substitute this back into the expression: $3+\frac{1}{3+\frac{1}{\frac{10}{3}}}=3+\frac{1}{3+\frac{3}{10}}.$ Next, simplify the fraction inside the denominator: $3+\frac{3}{10}=\frac{30}{10}+\frac{3}{10}=\frac{33}{10}.$ Now substitute this back into the expression: $3+\frac{1}{\frac{33}{10}}=3+\frac{10}{33}.$ To add these, we need a common denominator: $3+\frac{10}{33}=\frac{99}{33}+\frac{10}{33}=\frac{109}{33}.$ The fraction $\frac{109}{33}$ is already in its simplest form because 109 and 33 have no common factors other than 1. Therefore, $m=109$ and $n=33$, and the value of $m+n$ is: $109+33=142.$ So, the final answer is: $\boxed{142}.$
Co-rewarding-II Answer: To solve the given expression $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$, we need to simplify it step by step. First, simplify the innermost fraction: $3+\frac{1}{3}=\frac{9}{3}+\frac{1}{3}=\frac{10}{3}.$ Next, substitute this back into the next fraction: $3+\frac{1}{3+\frac{1}{\frac{10}{3}}}=3+\frac{1}{3+\frac{3}{10}}=3+\frac{1}{\frac{30}{10}+\frac{3}{10}}=3+\frac{1}{\frac{33}{10}}=3+\frac{10}{33}.$ Now, simplify the expression $3+\frac{10}{33}$: $3+\frac{10}{33}=\frac{99}{33}+\frac{10}{33}=\frac{109}{33}.$ The fraction $\frac{109}{33}$ is already in its simplest form because 109 and $3$ are coprime (they have no common factors other than 1). Thus, the irreducible fraction value of the given expression is $\frac{109}{33}$, where $m=109$ and $n=33$. Therefore, $m+n=109+33=142$. The value of $m+n$ is $\boxed{142}$.
GT-Reward Answer: Alright, I have this problem: The expression is: $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$ And I need to find the irreducible fraction $\frac{m}{n}$ that equals this expression, then determine $m+n$. First, I need to simplify the expression step by step. It looks like a continued fraction, so I’$l$ start by simplifying the innermost part and work my way out. Let’s denote the entire expression as: $E=3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$ **Step 1: Simplify the innermost fraction** Start with the innermost fraction: $3+\frac{1}{3}$ Calculate this: $3+\frac{1}{3}=\frac{9}{3}+\frac{1}{3}=\frac{10}{3}$ So now, the expression becomes: $E=3+\frac{1}{3+\frac{1}{\frac{10}{3}}}$ Wait, I need to simplify $\frac{1}{\frac{10}{3}}$: $\frac{1}{\frac{10}{3}}=\frac{3}{10}$ Now, the expression is: $E=3+\frac{1}{3+\frac{3}{10}}$ **Step 2: Simplify the next fraction** Now, simplify $\frac{3}{10}: 3+3$$\frac{3}{10}=\frac{30}{10}+\frac{3}{10}=\frac{33}{10}$ So, the expression now is: $E=3+\frac{1}{\frac{33}{10}}$ Simplify $\frac{1}{\frac{33}{10}}$: $\frac{1}{\frac{33}{10}}=\frac{10}{33}$ Now, the expression is: $E=3+\frac{10}{33}$ **Step 3: Combine the terms** Convert 3 to a fraction with denominator 33: $3=\frac{99}{33}$ Now, add the fractions: $E=\frac{99}{33}+\frac{10}{33}=\frac{109}{33}$ So, the expression simplifies to $\frac{109}{33}$. **Step 4: Ensure the fraction is irreducible** Check if 109 and 33 have any common divisors. 109 is a prime number (I think), and $3$ is $3\times 11$. Since 109 isn’t divisible by 3 or 11, the fraction $\frac{109}{33}$ is already in its irreducible form. **Step 5: Find m+nm+n** Here, $m=109$ and $n=33$, so: $m+n=109+33=142$ Wait, let’s double-check the calculations to ensure no mistakes were made. **Verification:** Starting over: $E=3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$ Innermost: $3+\frac{1}{3}=\frac{10}{3}$ Next layer: $3+\frac{1}{\frac{10}{3}}=3+\frac{3}{10}=\frac{33}{10}$ Outermost: $3+\frac{1}{\frac{33}{10}}=3+\frac{10}{33}=\frac{99}{33}+\frac{10}{33}=\frac{109}{33}$ Yes, it checks out. Therefore, $m+n=109+33=142$. $\boxed{142}$
Majority-Voting Answer: To find the value of $\frac{m}{n}$ as the irreducible fraction of the expression $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$, we need to simplify the expression step by step. First, let’s simplify the innermost fraction: $3+\frac{1}{3}=3+\frac{1}{3}=\frac{9}{3}+\frac{1}{3}=\frac{10}{3}$ Now, substitute this back into the expression: $3+\frac{1}{3+\frac{1}{\frac{10}{3}}}=3+\frac{1}{3+\frac{3}{10}}$ Next, simplify the fraction $\frac{3}{10}: 3+3$$\frac{3}{10}=\frac{30}{10}+\frac{3}{10}=\frac{33}{10}$ Now, substitute this back into the expression: $3+\frac{1}{\frac{33}{10}}=3+\frac{10}{33}$ To add these, we need a common denominator: $3+\frac{10}{33}=\frac{99}{33}+\frac{10}{33}=\frac{109}{33}$ So, the irreducible fraction value of the expression is $\frac{109}{33}$. Here, $m=109$ and $n=33$. Therefore, the value of $m+n$ is: $m+n=109+33=142$ Thus, the value of $m+n$ is $\boxed{0}$.
Self-Certainty Answer: Alright, I have this math problem to solve: **Problem:** I need to solve a math problem. The problem is about understanding how to solve math problems. I need to understand how to solve this problem so I can understand more about math. ### Understanding the Problem First, I need to understand what the problem is. I need to understand math so I can understand more about numbers and how to solve problems with numbers. Understanding math will help me understand more about how to understand problems. ### Understanding Math Problems After understanding the problem, I need to understand more about math problems. Understanding math problems will help me understand how to understand these problems. Understanding these problems will help me understand more about how to understand math. ### Understanding Mathematical Concepts After understanding math problems, I need to understand more about mathematical concepts. Understanding mathematical concepts will help me understand more about math. Understanding these concepts will help me understand more about how math helps me understand numbers. ### Understanding Mathematical Operations After understanding mathematical concepts, I need to understand more about mathematical operations. Understanding mathematical operations will help me understand more about how to understand these operations. Understanding these operations will help me understand more about how these operations help me understand math. ### Understanding Mathematical Operations in Math After understanding mathematical operations, I need to understand …… ### Understanding Mathematical Methods After understanding mathematical operations, I need to understand more about mathematical methods. …… ### Understanding How Mathematical Methods Help …… ### Understanding How These Methods Help Me Understand Math After understanding more about how these methods help me, I need to understand …… ### Understanding How These Methods Help Me Understand Math Problems After understanding how these methods help me understand math, I need to understand …… ### Understanding How These Methods Help Me Solve Math Problems After understanding how these methods help me understand these math problems, I need to understand …… ### Understanding How These Methods Help Me Solve These Problems After understanding how these methods help me solve these problems, I need to understand more about ……
Entropy Answer: To find the value of $\frac{m}{n}$ as the irreducible fraction of the expression: $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}$ Let’s simplify the expression step by step. 1. **Simplify the innermost fraction:** $3+\frac{1}{3}$ $3+\frac{1}{3}=\frac{3\times 3}{3}+\frac{1}{3}=\frac{9}{3}+\frac{1}{3}=\frac{9+1}{3}=\frac{10}{3}$ 2. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3}}=3+\frac{1}{\frac{10}{3}}$ $3+\frac{1}{\frac{10}{3}}=3+\frac{1\times 3}{10}=3+\frac{3}{10}=\frac{3\times 10}{10}+\frac{3}{10}=\frac{30}{10}+\frac{3}{10}=\frac{30+3}{10}=\frac{33}{10}$ 3. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}=3+\frac{1}{\frac{33}{10}}$ $3+\frac{1}{\frac{33}{10}}=3+\frac{1\times 10}{33}=3+\frac{10}{33}=\frac{3\times 33}{33}+\frac{10}{33}=\frac{99}{33}+\frac{10}{33}=\frac{99+10}{33}=\frac{109}{33}$ 4. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}=3+\frac{1}{\frac{109}{33}}$ $3+\frac{1}{\frac{109}{33}}=3+\frac{1\times 33}{109}=3+\frac{33}{109}=\frac{3\times 109}{109}+\frac{33}{109}=\frac{327}{109}+\frac{33}{109}=\frac{327+33}{109}=\frac{360}{109}$ 5. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}=3+\frac{1}{\frac{360}{109}}$ $3+\frac{1}{\frac{360}{109}}=3+\frac{1\times 109}{360}=3+\frac{109}{360}=\frac{3\times 360}{360}+\frac{109}{360}=\frac{1080}{360}+\frac{109}{360}=\frac{1080+109}{360}=\frac{1189}{360}$ 6. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}=3+\frac{1}{\frac{1189}{360}}$ …… 7. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}=3+\frac{1}{\frac{3927}{1189}}$ …… 8. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}}=3+\frac{1}{\frac{12970}{3927}}$ …… 9. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}}}=3+\frac{1}{\frac{42837}{12970}}$ …… 10. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}}}}=3+\frac{1}{\frac{141481}{42837}}$ …… 11. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}}}}}=3+\frac{1}{\frac{467280}{141481}}$ …… 12. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}}}}}}=3+\frac{1}{\frac{1543321}{467280}}$ …… 13. **Simplify the next fraction:** $3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3+\frac{1}{3}}}}}}}}}}}}}=3+\frac{1}{\frac{5097243}{1543321}}$ $3+\frac{1}{\frac{5097243}{1543321}}=3+$